OpenStudy (anonymous):

5. Determine whether the following two lines are parallel, perpendicular, neither, or both. L1: passes through (6, 4) and (3, 6) L2: passes through (5, 4) and (3, 1) both parallel and perpendicular perpendicular parallel neither parallel or perpendicular

OpenStudy (angela210793):

two parallel lines must have the same coefficient k1=(y2-y1)/(x2-x1)=(6-4)/(3-6)=-2/3 k2=(1-4)/(3-5)=3/2 so the lines aren't parallel 2 lines are perpendicular when K1*k2=-1---> -2/3*3/2=-1 so the lines are perpendicular :)

OpenStudy (angela210793):

(I'm not sure with wht you express the coefficient..but in Albania we use K that's why I used it :)

OpenStudy (anonymous):

oh you're an albanian ?

OpenStudy (anonymous):

the coefficient told by angela is also called the slope of the line...

OpenStudy (angela210793):

Yea :) Have you heard abt it?

OpenStudy (anonymous):

im KOSOVAR :)

OpenStudy (angela210793):

Woaaaaaaaaaaaawwwwwwwwwwwwwwwwwwww me te vertete?????????

OpenStudy (anonymous):

me te vertete he, hyn ne chat

OpenStudy (angela210793):

Are you kidding me?